a contractor mixes concrete from bags of pre-mix for small jobs. How many bags with 11% cement should he mix with 10 bags of 12.5% cement to produce a mix containing 12% cement?

Found 2 solutions by checkley77, nerdybill:
Answer by checkley77(12569)

.11x+.125*10=.12(10+x)
.11x+1.25=1.2+.12x
.11x-.12x=1.2-1.25
-.01x=-.05
x=-.05/-.01
x=5 bags of 12% is used.
Proof:
.11*5+.125*10=.12(10+5)
.55+1.25=.12*15
1.8=1.8

Answer by nerdybill(6960)

a contractor mixes concrete from bags of pre-mix for small jobs. How many bags with 11% cement should he mix with 10 bags of 12.5% cement to produce a mix containing 12% cement?
.
Let x = bags of 11% cement to be added
.
"cement from 11%" + "cement from 12.5%" = "total cement in mix"
.11x + .125(10) = .12(10+x)
.11x + 1.25 = 1.2 + .12x
1.25 = 1.2 + .01x
.05 = .01x
5 bags = x